# EEG2GAIT: A Hierarchical Graph Convolutional Network for EEG-based Gait Decoding

### v1

@@ Abstract
Decoding gait dynamics from EEG signals presents significant challenges due to the complex spatial dependencies of motor processes, the need for accurate temporal and spectral feature extraction, and the scarcity of high-quality gait EEG datasets. To address these issues, we propose EEG2GAIT, a novel hierarchical graph-based model that captures multi-level spatial embeddings of EEG channels using a Hierarchical Graph Convolutional Network (GCN) Pyramid. To further improve decoding accuracy, we introduce a Hybrid Temporal-Spectral Reward (HTSR) loss function, which combines time-domain, frequency-domain, and reward-based loss components. Moreover, we contribute a new Gait-EEG Dataset (GED), consisting of synchronized EEG and lower-limb joint angle data collected from 50 participants over two lab visits. Validation experiments on both the GED and the publicly available Mobile Brain-body imaging (MoBI) dataset demonstrate that EEG2GAIT outperforms state-of-the-art methods and achieves the best joint angle prediction. Ablation studies validate the contributions of the hierarchical GCN modules and HTSR Loss, while saliency maps reveal the significance of motor-related brain regions in decoding tasks. These findings underscore EEG2GAIT’s potential for advancing brain-computer interface applications, particularly in lower-limb rehabilitation and assistive technologies.
@@ 1 Introduction
Gait represents a fundamental aspect of human motor behavior, relying on motor control systems as well as cognitive and sensory processes[schmidt2018motor]. The EEG, which records electrical activity from the surface of the brain, provides a non-invasive means of monitoring these neural dynamics in real-time. Thus, decoding gait patterns from the electroencephalogram (EEG) constitutes a critical step in unraveling the complex interactions between brain activity and motor functions and offers valuable insights into the neural mechanisms that control movement[pfurtscheller1977event]. This capability proves especially crucial in neural rehabilitation, where accurate decoding of movement-related brain signals allows patients to control external devices with greater precision[wang2012self]. Accurate gait decoding furthermore promotes new therapeutic approaches within various neurological conditions such as stroke[baniqued2021brain], spinal cord injury[cui2022bci], and Parkinson’s disease[romero2024clinical].
@@
To achieve effective EEG-based motor decoding, various machine learning algorithms have been developed to date[ang2008filter, higashi2013common, costa2016decoding, miao2017spatial, riquelme2020better]. However, the mentioned works largely focused on manual feature extraction from the EEG, failing to extract highly discriminative features unique to each participant and thus limiting the decoding performance. Deep learning algorithms have addressed these limitations to some extent by extracting subject-specific EEG patterns in a data-driven manner, thereby enhancing the decoding performance. For instance, long short-term memory (LSTM) networks[wang2018lstm] have been utilized to capture the temporal dependencies in EEG signals, reflecting the dynamic nature of brain activity during gait cycles. Convolutional neural networks (CNNs) have likewise been employed to automatically learn spatial-temporal features from raw EEG signals, improving the identification of movement-related patterns[schirrmeister2017deep, lawhern2018eegnet]. These advanced models leverage the intricate relationship between neural activations and motor execution, enabling more precise and robust decoding of gait patterns.
@@
Building upon this foundation, neurokinematics research has further highlighted the coordinated activation of different brain regions during various stages of movement, each contributing to distinct cognitive processes[reddy2018brain]. Numerous studies have explored brain region activations during motor tasks using EEG, providing valuable insights into the underlying neurophysiological processes. Event-related desynchronization (ERD) and synchronization (ERS) have been widely used to characterize motor cortex activity during movement planning and execution[pfurtscheller1999event]. For instance, $\mu$ and $\beta$ rhythms, which are associated with motor actions, show distinctive patterns of modulation during motor tasks, highlighting the dynamic activation of cortical regions[neuper2001event]. Additionally, EEG changes observed during motor skill acquisition, particularly in the $\alpha$ and $\beta$ bands, reflect cortical plasticity and the reorganization of motor areas as learning progresses, a process influenced by both age and task complexity[bootsma2021neural]. The negative correlation between EEG $\alpha$/$\beta$-band activity and BOLD signals in sensorimotor areas during motor tasks has been shown to provide valuable insights into the dynamics of motor cortex engagement[yuan2010negative]. The use of EEG to reconstruct three-dimensional hand movements has further demonstrated its efficacy in real-time decoding of motor region activity[bradberry2010reconstructing]. Inspired by these findings, several deep learning approaches have incorporated prior neurophysiologal knowledge to improve motor decoding performance. Functional topological relationships between electrodes are leveraged to decode time-resolved EEG motor imagery signals using Graph Convolutional Networks (GCNs), as demonstrated in prior work [hou2022gcns]. Local and global biological topology among brain regions has also been utilized for EEG-based emotion recognition through graph neural networks with regularization, as shown by Zhong and colleagues [zhong2020eeg]. These graph-based methods highlight the benefits of treating EEG electrodes as a non-Euclidean structure, enabling enhanced motor decoding performance.
@@
To address the shortcomings of existing models, we propose the new model EEG2GAIT which introduces a Hierarchical GCN Pyramid for EEG-based gait regression. This model is specifically designed to capture the intricate spatial relationships among EEG channels by extracting multi-level spatial embeddings that effectively represent the dynamic brain activity involved in gait.
To further improve decoding accuracy, we propose a novel loss function, the Hybrid Temporal-Spectral Reward (HTSR) Loss, which is specifically designed for time series modelling. This loss function integrates three essential components: an MSE Loss, a Time-Frequency Loss, and a Reward Loss. The MSE Loss serves as the core metric, ensuring that the predicted gait values align closely with the actual data in the time domain. The time-frequency component enhances this by capturing both time and frequency characteristics of the gait pattern, enabling the model to account for long-term dependencies and dynamic patterns often overlooked by conventional time-series models. Additionally, the Reward Loss encourages the model to give due importance to well-predicted samples, preventing underutilization of global features and ensuring that the model learns from both large and small errors. By addressing the subtle variations in gait patterns, HTSR Loss improves the model’s ability to generalize across different datasets, offering a robust and comprehensive solution for continuous gait decoding from EEG signals.
Additionally, to enrich the existing frame of high-quality gait datasets, we contribute a new walking dataset consisting of EEG and joint angle data of 50 able-bodied participants. This dataset includes natural gait data collected from each participant over two visits to the lab on different days.
We present experimental results and statistical analyses comparing EEG2GAIT with state-of-art baseline methods. Experimental results demonstrate that the proposed method achieves the highest regression performance among the compared methods. Furthermore, we conducted ablation studies to demonstrate the contribution of each module introduced within this work. In addition, we generated saliency maps to visualize spatial correlations and gain insights into the features learned by the model.
@@
  • We propose EEG2GAIT Model: We propose a novel model, EEG2GAIT, incorporating a Hierarchical GCN Pyramid for EEG-based gait regression. The model effectively captures complex spatial relationships in EEG data.
  • We propose Hybrid Temporal-Spectral Reward (HTSR) Loss: We introduce a new loss function, HTSR Loss, designed for time series modelling which combines time-, time-frequency-, and reward-based losses.
  • We collect a New Walking Dataset: In addition to these methodological advancements, we collect a novel dataset consisting of gait data from 50 participants across two lab visits. We conduct comprehensive experiments, including ablation studies and performance evaluation, to assess the effectiveness of HTSR loss and EEG2GAIT. The results demonstrate that EEG2GAIT outperforms all state-of-the-art models, while HTSR loss significantly enhances the model’s performance, leading to improved prediction accuracy and robustness.
@@ 2.1 Neural Decoding of Motor Execution Using EEG Signals
EEG-based neural decoding of lower-limb motor execution has been widely studied, leveraging EEG’s high temporal resolution for capturing motor control dynamics. Pfurtscheller and colleagues [pfurtscheller2001motor] demonstrated that event-related desynchronization (ERD) and synchronization (ERS) effectively characterize motor execution and imagery, laying the groundwork for decoding lower-limb movements. Ang and colleagues previously [ang2012filter] used common spatial pattern (CSP) and linear discriminant analysis (LDA) to decode gait phases, enhancing motor-related EEG feature extraction. Schirrmeister and colleagues [schirrmeister2017deep] employed convolutional neural networks (CNNs), boosting accuracy in classifying lower-limb motor imagery. Recent deep learning methods have improved feature extraction and spatial representation of EEG signals[goh2018spatio, tortora2020deep, fu2022matn], demonstrating their capability to capture complex EEG patterns. Finally, Wang and colleagues[wang2018implementation] developed a real-time BCI system for lower-limb exoskeleton control, demonstrating EEG’s feasibility in rehabilitation. These studies have contributed to the development of BCIs for rehabilitation and assistive technologies.
@@ 2.2 Graph Neural Networks
Graph Neural Networks (GNNs) are designed for processing non-Euclidean data with a graph-like structure. One category of GNNs, known as spectral GNNs, often depends on costly eigendecomposition of the graph Laplacian[bruna2013spectral]. To address this limitation, several techniques use approximation methods for spectral filtering. For instance, ChebyNet[defferrard2016convolutional] leverages Chebyshev polynomials to approximate spectral filters, while Cayley polynomials are employed to compute spectral filters targeting specific frequency bands, scaling linearly for sparse graphs[levie2018cayleynets]. GCN simplifies spectral filtering through localized first-order aggregation[kipf2022semi]. Since EEG signals naturally exhibit a graph structure, some approaches apply GNNs/GCNs to capture spatial features from EEG data. Several methods [shamsi2021early, hou2022gcns, zhang2020motor] model brain activity using a single adjacency matrix, which provides a global representation but may not fully capture complex inter-regional interactions. Furthermore, some approaches [hou2022gcns, shamsi2021early, zhang2020motor] focus on spatial graph structures, often using averaged node features, which simplifies modeling but may reduce temporal specificity.
@@ 2.3 Enhancing Model Performance with Well-Classified Samples
In regression tasks, MSE, MAE, and Root Mean Squared Error (RMSE) are commonly used as loss functions. However, these losses prioritize poorly classified samples, as well-classified ones offer limited new information for the model. Recent studies have shown that leveraging well-classified samples can enhance model performance. For example, [zhao2022well] introduces a loss function that prioritizes well-classified samples, thus enhancing feature extraction and classification accuracy. In DeepNoise[yang2022deepnoise], well-classified samples were used to improve accuracy in noisy fluorescence microscopy images. Jung and colleagues [jung2022wbc] used generative models for white blood cell classification, demonstrating that high-quality samples improve model performance. Additionally, in EEG signal decoding, Kwon and colleagues [kwon2022friend] propose a strategy that effectively leverages well-classified samples by introducing carefully calculated friend and enemy noise to retain correct classification in friendly models while inducing misclassification in adversarial ones. These studies suggest that well-classified samples can enhance model robustness and generalization under challenging conditions.
3 Methods
@@ 3.1 EEG2GAIT Overview
Figure 1: Overview of EEG2GAIT Architecture and Loss Calculation. The sequential EEG signals are segmented into training samples $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time length. The Local Temporal Learner (LTL) uses 1D convolutions along the temporal dimension to capture local temporal information. The output is then reorganised into $n\times T$ temporal graphs, where $n$ is the number of samples. These graphs are input to the Hierarchical GCN Pyramid (HGP), which transforms each graph into a token embedding. The token for all graphs and the input for graph construction will be accumulated as input to the Global Spatial Learner (GSL). GSL performs 1D convolutions across all channels. The outputs are passed into the Feature Fusion Layers and then Global Temporal Learner (GTL) for global temporal information extraction via self-attention. The final output goes through the Output Layer. Losses from both the temporal (MSE) and frequency (FFT-based L1) domains are combined via a Reward Loss block to yield the Hybrid Temporal-Spectral Reward Loss.
@@ 3.2 Local Temporal Learner
Let a time-series EEG signal be denoted as $(X,Y)$, where $X\in\mathbb{R}^{C\times t}$ represents the EEG signal sampled at 100 Hz, and $Y\in\mathbb{R}^{d_{J}\times t}$ denotes the corresponding label, with $C$ denoting the number of channels, $t$ the total number of recorded time points, and $d_{J}$ representing the number of lower-limb joint angles. Subsequently, a sliding window of length $T$ and step size 1 is applied to further segment the signal into shorter subsequences. As a result, $X$ is divided into $(t-T+1)$ samples, denoted individually as $\bar{X}\in\mathbb{R}^{C\times T}$. Each $\bar{X}$ corresponds to a label $\bar{y}\in\mathbb{R}^{d_{J}}$.
@@ 3.3 Graph Construction Module
To enhance the flexibility of the graph representation, the adjacency matrix is treated as a learnable parameter during training. This dynamic learning process allows the model to adaptively adjust the connections between channels, capturing task-specific spatial relationships inherent in the EEG signals. The updated adjacency matrix $\tilde{A}$ at any training step is computed as:
$\tilde{A}=\text{ReLU}(A+A^{\top})+I,$
where $I$ is the identity matrix representing self-loops, and ReLU ensures non-negative edge weights.

### v2

@@ Abstract
Decoding gait dynamics from EEG signals presents significant challenges due to the complex spatial dependencies of motor processes, the need for accurate temporal and spectral feature extraction, and the scarcity of high-quality gait EEG datasets. To address these issues, we propose EEG2GAIT, a novel hierarchical graph-based model that captures multi-level spatial embeddings of EEG channels using a Hierarchical Graph Convolutional Network (GCN) Pyramid. To further improve decoding performance, we introduce a Hybrid Temporal-Spectral Reward (HTSR) loss function, which integrates time-domain, frequency-domain, and reward-based loss components. In addition, we contribute a new Gait-EEG Dataset (GED), consisting of synchronized EEG and lower-limb joint angle data collected from 50 participants across two laboratory visits. Extensive experiments demonstrate that EEG2GAIT with HTSR achieves superior performance on the GED dataset, reaching a Pearson correlation coefficient ($r$) of 0.959, a coefficient of determination ($R^{2}$) of 0.914, and a Mean Absolute Error (MAE) of 0.193. On the MoBI dataset, EEG2GAIT likewise consistently outperforms existing methods, achieving an $r$ of 0.779, an $R^{2}$ of 0.597, and an MAE of 4.384. Statistical analyses confirm that these improvements are significant compared to all prior models. Ablation studies further validate the contributions of the hierarchical GCN modules and the proposed HTSR loss, while saliency analysis highlights the involvement of motor-related brain regions in decoding tasks. Collectively, these findings underscore EEG2GAIT’s potential for advancing brain-computer interface applications, particularly in lower-limb rehabilitation and assistive technologies.
@@ 1 Introduction
Gait is a fundamental component of human motor behavior, relying on not only motor control systems but also cognitive and sensory processes [39]. The electroencephalogram (EEG), which records electrical activity from the scalp, provides a non-invasive way to monitor these neural dynamics in real time. Thus, decoding gait patterns from EEG signals is a critical step toward understanding the complex interplay between brain activity and motor functions, offering valuable insights into the neural mechanisms underlying movement [32]. This capability is particularly crucial in neural rehabilitation, where accurate decoding of movement-related brain signals enables patients to control external devices with greater precision [50]. Moreover, accurate gait decoding facilitates the development of new therapeutic approaches for various neurological conditions, including stroke [3], spinal cord injury [9], and Parkinson’s disease [37].
@@
To achieve effective EEG-based motor decoding, various machine learning algorithms have been developed to date [2, 16, 8, 28, 36]. However, the mentioned works largely focused on manual feature extraction from the EEG, failing to extract highly discriminative features unique to each participant and thus limiting the decoding performance. Deep learning algorithms have addressed these limitations to some extent by extracting subject-specific EEG patterns in a data-driven manner, thereby enhancing the decoding performance. For instance, long short-term memory (LSTM) networks [49] have been utilized to capture the temporal dependencies in EEG signals, reflecting the dynamic nature of brain activity during gait cycles. Convolutional neural networks (CNNs) have likewise been employed to automatically learn spatial-temporal features from raw EEG signals, improving the identification of movement-related patterns [38, 25]. These advanced models leverage the intricate relationship between neural activations and motor execution, enabling more precise and robust decoding of gait patterns.
@@
Building upon this foundation, neurokinematics research has further highlighted the coordinated activation of different brain regions during various stages of movement, each contributing to distinct cognitive processes [35]. Numerous studies have explored brain region activations during motor tasks using EEG, providing valuable insights into the underlying neurophysiological processes. Event-related desynchronization (ERD) and synchronization (ERS) have been widely used to characterize motor cortex activity during movement planning and execution [33]. For instance, $\mu$ and $\beta$ rhythms, which are associated with motor actions, show distinctive patterns of modulation during motor tasks, highlighting the dynamic activation of cortical regions [30]. Additionally, EEG changes observed during motor skill acquisition, particularly in the $\alpha$ and $\beta$ bands, reflect cortical plasticity and the reorganization of motor areas as learning progresses, a process influenced by both age and task complexity [4]. The negative correlation between EEG $\alpha$$\beta$-band activity and BOLD signals in sensorimotor areas during motor tasks has been shown to provide valuable insights into the dynamics of motor cortex engagement [54]. The use of EEG to reconstruct three-dimensional hand movements has further demonstrated its efficacy in real-time decoding of motor region activity [5]. Inspired by these findings, several deep learning approaches have incorporated prior neurophysiologal knowledge to improve motor decoding performance. Functional topological relationships between electrodes are leveraged to decode time-resolved EEG motor imagery signals using Graph Convolutional Networks (GCNs), as demonstrated in prior work [17]. Local and global biological topology among brain regions has also been utilized for EEG-based emotion recognition through graph neural networks with regularization, as shown by Zhong and colleagues [58]. These graph-based methods highlight the benefits of treating EEG electrodes as a non-Euclidean structure, enabling enhanced motor decoding performance.
@@
To address the shortcomings of existing models, we propose the new model EEG2GAIT which introduces a Hierarchical GCN Pyramid (HGP) for EEG-based gait regression. This model is specifically designed to capture the intricate spatial relationships among EEG channels by extracting multi-level spatial embeddings that effectively represent the dynamic brain activity involved in gait.
To further improve decoding accuracy, we propose a novel loss function, the Hybrid Temporal-Spectral Reward (HTSR) Loss, specifically designed for time series modelling. This loss function integrates three essential components: an MSE Loss, a Time-Frequency Loss, and a Reward Loss. The MSE Loss serves as the core metric, ensuring that the predicted gait values align closely with the actual data in the time domain. The time-frequency component enhances this by capturing both temporal and spectral characteristics of gait patterns, enabling the model to account for long-term dependencies and dynamic patterns often overlooked by conventional time-series models. Additionally, the Reward Loss encourages the model to focus on well-predicted samples, preventing underutilization of global features and ensuring balanced learning from both large and small errors. By addressing subtle variations in gait patterns, HTSR Loss aims to improve model generalizability across different datasets, offering a robust and comprehensive approach for continuous gait decoding from EEG signals.
To enrich the existing body of high-quality gait datasets, we also contribute a new walking dataset comprising EEG and joint angle data from 50 able-bodied participants. The dataset includes natural gait data collected from each participant across two visits to the laboratory on different days.
In this work, we propose EEG2GAIT, a novel framework designed to decode gait patterns from EEG signals. We benchmark EEG2GAIT against several state-of-the-art baselines and conduct comprehensive ablation studies to analyze the contributions of individual modules. Additionally, we incorporate saliency map visualization to investigate the spatial characteristics of learned neural representations and to better understand the relationship between EEG features and motor behavior.
@@
  • EEG2GAIT Model: A novel model incorporating a Hierarchical Graph Convolutional Network Pyramid for EEG-based gait regression, effectively capturing complex spatial relationships in EEG data.
  • Hybrid Temporal-Spectral Reward (HTSR) Loss: A newly designed loss function for time series modelling that combines time-, time-frequency-, and reward-based objectives to enhance decoding accuracy and robustness.
  • New Walking Dataset: A large-scale dataset collected from 50 participants over two lab visits, supporting extensive evaluation and future research in EEG-based gait decoding.
@@ 2.1 Neural Decoding of Motor Execution Using EEG Signals
EEG-based neural decoding of lower-limb motor execution has been widely studied, leveraging EEG’s high temporal resolution for capturing motor control dynamics. Pfurtscheller and colleagues [34] demonstrated that event-related desynchronization (ERD) and synchronization (ERS) effectively characterize motor execution and imagery, laying the groundwork for decoding lower-limb movements. Ang and colleagues previously [1] used common spatial pattern (CSP) and linear discriminant analysis (LDA) to decode gait phases, enhancing motor-related EEG feature extraction. Schirrmeister and colleagues [38] employed convolutional neural networks (CNNs), boosting accuracy in classifying lower-limb motor imagery. Recent deep learning methods have improved feature extraction and spatial representation of EEG signals [14, 46, 12], demonstrating their capability to capture complex EEG patterns. Finally, Wang and colleagues [48] developed a real-time BCI system for lower-limb exoskeleton control, demonstrating EEG’s feasibility in rehabilitation. These studies have contributed to the development of BCIs for rehabilitation and assistive technologies.
@@ 2.2 Graph Neural Networks
Graph Neural Networks (GNNs) are designed for processing non-Euclidean data with a graph-like structure. One category of GNNs, known as spectral GNNs, often depends on costly eigendecomposition of the graph Laplacian [6]. To address this limitation, several techniques use approximation methods for spectral filtering. For instance, ChebyNet [10] leverages Chebyshev polynomials to approximate spectral filters, while Cayley polynomials are employed to compute spectral filters targeting specific frequency bands, scaling linearly for sparse graphs [26]. GCN simplifies spectral filtering through localized first-order aggregation [22]. Since EEG signals naturally exhibit a graph structure, some approaches apply GNNs/GCNs to capture spatial features from EEG data. Several methods [40, 17, 55] model brain activity using a single adjacency matrix, which provides a global representation but may not fully capture complex inter-regional interactions. Furthermore, some approaches [17, 40, 55] focus on spatial graph structures, often using averaged node features, which simplifies modeling but may reduce temporal specificity.
@@ 2.3 Enhancing Model Performance with Well-Classified Samples
In regression tasks, MSE, MAE, and Root Mean Squared Error (RMSE) are commonly used as loss functions. However, these losses prioritize poorly classified samples, as well-classified ones offer limited new information for the model. Recent studies have shown that leveraging well-classified samples can enhance model performance. For example, [57] introduces a loss function that prioritizes well-classified samples, thus enhancing feature extraction and classification accuracy. In DeepNoise [53], well-classified samples were used to improve accuracy in noisy fluorescence microscopy images. Jung and colleagues [21] used generative models for white blood cell classification, demonstrating that high-quality samples improve model performance. Additionally, in EEG signal decoding, Kwon and colleagues [24] propose a strategy that effectively leverages well-classified samples by introducing carefully calculated friend and enemy noise to retain correct classification in friendly models while inducing misclassification in adversarial ones. These studies suggest that well-classified samples can enhance model robustness and generalization under challenging conditions.
3 Methods and Materials
@@ 3.1 EEG2GAIT Overview
Figure 1: Overview of EEG2GAIT Architecture and Loss Calculation. The sequential EEG signals are segmented into training samples $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time window length. The Local Temporal Learner (LTL) uses 1D convolutions along the temporal dimension to capture local temporal information. Each EEG sequence is divided into $n$ segments, and the LTL output is then reorganized into $n$ temporal graphs, each representing one segment. These graphs are input to the Hierarchical GCN Pyramid (HGP), which transforms each graph into a token embedding. The token for all graphs and the input for graph construction will be accumulated as input to the Global Spatial Learner (GSL). GSL performs 1D convolutions across all channels. The outputs are passed into the Feature Fusion Layers and then Global Temporal Learner (GTL) for global temporal information extraction via self-attention. The final output goes through the Output Layer. Losses from both the temporal (MSE) and frequency (FFT-based L1) domains are combined via a Reward Loss block to yield the Hybrid Temporal-Spectral Reward Loss.
@@ 3.2 Local Temporal Learner
Let a time-series EEG signal be denoted as $(X,Y)$, where $X\in\mathbb{R}^{C\times t}$ represents the EEG signal sampled at 100 Hz, and $Y\in\mathbb{R}^{d_{J}\times t}$ denotes the corresponding label, with $C$ denoting the number of channels, $t$ the total number of recorded time points, and $d_{J}$ representing the number of lower-limb joint angles. Subsequently, a sliding window of length $T$ and step size 1 is applied to further segment the signal into shorter subsequences. As a result, $X$ is divided into (tT+1)(t-T+1) samples, denoted individually as $\bar{X}\in\mathbb{R}^{C\times T}$. Each $\bar{X}$ corresponds to a label $\bar{y}\in\mathbb{R}^{d_{J}}$.
@@ 3.3 Graph Construction Module
To enhance the flexibility of the graph representation, the adjacency matrix is implemented as a learnable parameter (nn.Parameter) during training. This dynamic parameterization allows the network to adaptively refine the connectivity structure rather than relying solely on fixed priors.
To provide a biologically meaningful initialization, we first construct an anatomical adjacency matrix $A_{\text{prior}}\in\mathbb{R}^{C\times C}$ based on spatial priors: two channels are connected if their inter-electrode distance is less than 30 mm, yielding a sparse, topology-informed graph. The prior is pre-processed as:
$\color[rgb]{0,0,0}\tilde{A}_{\text{prior}}=\text{ReLU}(A_{\text{prior}}+A_{\text{prior}}^{\top})+I,$} (1)
where $I$ denotes the identity matrix (self-loops), and ReLU guarantees non-negative edge weights.
During training, the learnable adjacency parameter $A$ (initialized from $\tilde{A}_{\text{prior}}$) is dynamically normalized in each forward pass using symmetric normalization:
$D$ $=\text{diag}\left(\sum_{j}A_{ij}+\text{mask}_{i}\right)$ (2)
$\text{mask}_{i}$ $=\begin{cases}1&\text{if }\sum_{j}A_{ij}=0\\ 0&\text{otherwise}\end{cases}$ (3)
$\hat{A}$ $=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ (4)
where the mask ensures numerical stability for isolated nodes, and the symmetric normalization $\hat{A}$ preserves the scale of feature representations during graph convolution.